# Time-Aware One Step Diffusion Network for Real-World Image Super-Resolution

### v2

@@ Abstract
Diffusion-based real-world image super-resolution (Real-ISR) methods have demonstrated impressive performance. To achieve efficient Real-ISR, many works employ Variational Score Distillation (VSD) to distill pre-trained stable-diffusion (SD) model for one-step SR with a fixed timestep. However, due to the different noise injection timesteps, the SD will perform different generative priors. Therefore, a fixed timestep is difficult for these methods to fully leverage the generative priors in SD, leading to suboptimal performance. To address this, we propose a Time-Aware one-step Diffusion Network for Real-ISR (TADSR). We first introduce a Time-Aware VAE Encoder, which projects the same image into different latent features based on timesteps. Through joint dynamic variation of timesteps and latent features, the student model can better align with the input pattern distribution of the pre-trained SD, thereby enabling more effective utilization of SD’s generative capabilities. To better activate the generative prior of SD at different timesteps, we propose a Time-Aware VSD loss that bridges the timesteps of the student model and those of the teacher model, thereby producing more consistent generative prior guidance conditioned on timesteps. Additionally, though utilizing the generative prior in SD at different timesteps, our method can naturally achieve controllable trade-offs between fidelity and realism by changing the timestep condition. Experimental results demonstrate that our method achieves both state-of-the-art performance and controllable SR results with only a single step.
Introduction
Image Super-Resolution (ISR) aims to reconstruct a high-quality (HQ) image from its low-quality (LQ) counterpart. Different from the simple degradation assumptions in traditional ISR (Lim et al. 2017; Zhang et al. 2018b; Chen et al. 2021; Liang et al. 2021; Chen et al. 2023; Dong et al. 2014), Real-World Image Super-Resolution (Real-ISR) aims to restore HQ images from LQ inputs degraded by complex and unknown factors in real-world scenarios, which has recently attracted increasing attention (Wang et al. 2021; Xie et al. 2023; Zhang et al. 2021; Liang, Zeng, and Zhang 2022).
To address a broader spectrum of degradation types and achieve more realistic results, many researchers have turned to generative models. Recently, diffusion models (Ho, Jain, and Abbeel 2020) have demonstrated a superior ability to generate fine-grained details in image generation tasks. Consequently, several works have explored leveraging the generative priors in pre-trained Stable Diffusion (SD) models (Rombach et al. 2022) to tackle Real-ISR, yielding impressive results (Wang et al. 2023a; Lin et al. 2023; Tao et al. 2023; Wu et al. 2024c; Duan et al. 2025). Nevertheless, the iterative denoising process inherent in diffusion models introduces significant computational overhead and latency.
@@ Introduction
Figure 1: (a) Comparison between our TADSR and PisaSR. In PisaSR, increasing the semantic weight $\lambda_{sem}$ leads to the restoration of more realistic images. As the timestep condition $t_{s}$ increases, our model demonstrates a significant improvement in generative capability, recovering a more realistic parrot image. In contrast, PisaSR shows only an increase in sharpness as $\lambda_{sem}$ increases. (b) The input image and the corresponding outputs of the teacher model at different timesteps $t$ with the VSD loss. The outputs of the teacher model vary significantly across different timesteps, reflecting distinct guidance orientations.
@@
To overcome these limitations, some researchers have focused on distilling SD into an efficient one-step model for Real-ISR (Wu et al. 2024b; Sun et al. 2025; Chen et al. 2025; Dong et al. 2025; Zhang et al. 2024). Specifically, OSEDiff (Wu et al. 2024b) first leverages the Variational Score Distillation (VSD) loss (Wang et al. 2023b) to distill the SD model, enabling realistic image reconstruction in a single step. Subsequently, S3Diff (Zhang et al. 2024), PisaSR (Sun et al. 2025), AdcSR (Chen et al. 2025), and TSDSR (Dong et al. 2025) also adopt distillation-based approaches to develop SD-based one-step Real-ISR models, using either adversarial loss or modified VSD loss.
@@
However, these works generally fix the injected timestep, e.g. step 999, and use the original VAE encoder in SD, which prevent them from effectively leveraging the generative prior in SD. Specifically, as shown in Fig. 1(b), the output of the teacher model (pre-trained SD) varies depending on the input timestep and latent features. When timestep $t$ equals 100, there are only differences in texture details between the output of the teacher model and the input image. In contrast, when $t$ increases to 300, the output shows clear differences, reflecting more of the leaf-related semantic prior learned by SD. However, with $t$ increasing further to 600, most of the image information is lost, and the teacher model can only recover the overall structure and color of the leaves, failing to provide meaningful guidance. These observations imply that fixed timestep and latent features are insufficient to fully activate the generative priors across different timesteps in SD, the gradient guidance from the teacher model is also timestep-dependent. As a result, as shown in Fig. 1(a), although we increase the semantic weight $\lambda_{sem}$ in PisaSR (Sun et al. 2025), it only produces enhanced sharpness without significantly enriching the semantic content.
@@
In this paper, we propose Time-Aware One Step Diffusion Network for Super-Resolution (TADSR), a framework that more effectively distills the generative prior of SD at different timesteps into a one-step diffusion model for Real-ISR. Our TADSR consists of two key components: (1) Time-Aware VAE Encoder (TAE): we introduce a time embedding layer into the VAE encoder, enabling it to map the same image to different latent representations based on the timestep, thereby achieving coordinated adjustment between the timestep and latent representation. Through the TAE, we can establish a matching relationship between the input timestep and latent representation similar to that in the original SD, allowing for more effective utilization of the generative priors of SD at different timesteps. (2) Time-Aware Variational Score Distillation (TAVSD) Loss: we design a mapping function to associate the timestep injected into the SR network with the one used in the VSD loss. When the SR network is conditioned on a larger timestep, the teacher receives a latent image corrupted with stronger noise, providing guidance that emphasizes stronger semantic generation in the reconstruction results. Conversely, smaller timesteps lead to similar results with reconstruction, primarily enhancing texture details. Therefore, TAVSD can provide more consistent generative guidance condition on the injected timestep in the SR network.
Thanks to TAE and TAVSD, the proposed TADSR can naturally leverage the generative priors of SD at different timesteps to achieve controllable trade-offs between fidelity and realism in Real-ISR and superior performance. As shown in Fig. 1(a), our method gradually generates a more realistic parrot as the timestep increases, though fully leveraging the generative priors in SD.
Related Work
Real-World Image Super-Resolution
@@
Traditional ISR methods (Lim et al. 2017; Zhang et al. 2018b; Chen et al. 2021; Liang et al. 2021; Chen et al. 2023; Dong et al. 2014; Zhao et al. 2025) typically degrade HQ images using simple downsampling operations to construct HQ-LQ image pairs for training, supervised by pixel-level losses. However, these approaches struggle to handle images degraded by complex real-world processes and often result in overly smooth reconstructions. To better simulate the unknown and complex degradations in real-world scenarios, several studies (Zhang et al. 2021; Wang et al. 2021) have proposed more sophisticated degradation pipelines to synthesize LQ data. Specifically, BSRGAN (Zhang et al. 2021) introduces a random combination of basic degradation operations (e.g., downsampling, blurring, noise) injection, with varying intensities to generate more realistic HQ-LQ pairs. Real-ESRGAN (Wang et al. 2021) proposes a second-order degradation scheme to cover a broader range of degradation types. In addition, inspired by Generative Adversarial Networks (GANs), researchers have incorporated discriminators into Real-ISR frameworks and adopted adversarial losses to encourage the reconstruction of more realistic images. Although these GAN-based methods can produce richer texture details compared to traditional approaches, they are often unstable to train and prone to generating unnatural artifacts (Wu et al. 2024b).
@@
Diffusion-Based Real-ISR
Recently, many researchers have leveraged the powerful generative priors of pre-trained diffusion models for Real-ISR tasks to achieve high-fidelity image reconstruction. For example, StableSR (Wang et al. 2023a) conditions the diffusion process on LQ images by injecting them through a learnable time-aware encoder into the SD model, enabling strong detail generation capabilities. DiffBIR (Lin et al. 2023) utilizes ControlNet to extract structural information from LQ images and incorporates it as a control signal to better guide the generative prior of SD for super-resolution. PASD (Tao et al. 2023) and SeeSR (Wu et al. 2024c) extract semantic information from LQ inputs and inject it into SD, resulting in more realistic super-resolved outputs.
Although these approaches yield impressive results, the multi-step denoising process leads to high computational and time costs. To accelerate diffusion-based Real-ISR, OSEDiff introduces the VSD loss to distill the pre-trained SD model, enabling realistic image reconstruction in a single step by only training LoRA parameters mounted on SD. S3Diff further adopts degradation-guided LoRA adapters combined with adversarial training to achieve one-step super-resolution. PisaSR trains two separate LoRA adapters for pixel-level and semantic-level guidance, allowing controllable trade-offs between realism and fidelity.
However, these methods overlook the varying generative capabilities of SD in different timesteps: larger timesteps favor semantic and structural generation, while smaller timesteps emphasize texture and detail. Our work aims to fully exploit these time-dependent generation characteristics to achieve superior SR performance and a natural balance between visual fidelity and generative realism.
Methodology
Problem Definition
Real-ISR aims to reconstruct HQ images $x_{H}$ from LQ images $x_{L}$ that suffer from complex and unknown degradations. With the advancement of deep learning, researchers have commonly adopted neural networks $G_{\theta}$ to estimate the HQ images and optimize the network through loss functions. The general form of the loss function is as:
@@
Recently, with the advancement of diffusion models, several studies (Wu et al. 2024b; Sun et al. 2025) have leveraged the generative prior in pre-trained SD and adopted VSD as a regression objective. The formation of VSD is:
@@
where $\epsilon_{\psi}$ is the pre-trained diffusion model (teacher model), $\epsilon_{\phi}$ represents its replica with trainable LoRA (LoRA model), and $c$ is a text embedding of a caption describing the input image. $\hat{z}=G_{\theta}(x_{L})$ is the output of the student network $G_{\theta}$, and $\hat{z}_{t}=\alpha_{t}\hat{z}+\beta_{t}\epsilon$ is the noisy latent input. $\epsilon$ is the gaussian noise, and $\alpha_{t}$ and $\beta_{t}$ are the scale parameters in diffusion.
Formally, the VSD loss can be viewed as the residual between the noise outputs of the teacher model and the LoRA model, which is equivalent to the residual of their predicted latent images (detailed proof can be found in supplementary material). Therefore, in the following, we represent the VSD loss using the latent images and their residual.
Overview
As illustrated in Fig. 2, we distill a Student Model $G_{\theta}$ to perform one-step Real-ISR, which is consist of a Time-Aware Encoder $E_{\theta}$ and a UNet $F_{\theta}$. First, we sample an HQ-LQ image pair from the dataset and a timestep $t_{s}$. Then, both the LQ image and $t_{s}$ are fed into Student Model $G_{\theta}$ to obtain the latent output $\hat{z}_{0}$. We decode $\hat{z}_{0}$ into pixel space and compute the reconstruction loss with the HQ image to ensure fidelity. In the latent space, we map the $t_{s}$ to another timestep $t_{v}$, and add noise corresponding to timestep $t_{v}$ to the $\hat{z}_{0}$ to obtain $\hat{z}_{t_{v}}$. Then, we feed $\hat{z}_{t_{v}}$ and $t_{v}$ into both the teacher model and the LoRA model to compute the Time-Aware Variational Score Distillation loss $\mathcal{L}_{TAVSD}$ to enhance the realism. Consistent with Fig. 1(b), when $t_{v}$ is small, the gradients produced by the TAVSD loss are relatively small and mainly reflect texture details. In contrast, when $t_{v}$ is large, the gradients become significantly large and provide more semantic guidance.
@@
Time-Aware VAE Encoder
In the original SD, as timestep increases, the latent features fed into the model are injected with more noise, thereby activating different generative priors to produce the image. However, current one-step SD-based Real-ISR methods typically adopt a fixed timestep during training, making it difficult to fully exploit the generative priors in SD. Since the original VAE encoder can only encode the same image into a single latent feature, simply sampling the timestep randomly is insufficient, which is clearly inconsistent with the correlation between timesteps and latent distributions in SD.
To better utilize the generative priors in SD, we propose a Time-Aware VAE Encoder (TAE). By incorporating a temporal embedding layer into the VAE encoder, TAE encodes the input image into different latent distributions conditioned on the timestep $t_{s}$, enabling synchronized variation between $t_{s}$ and latent distribution, thus better activating the generative priors at different timesteps within SD. This process can be formulated as:

### v3

# arXiv:2508.16557v3[eess.IV] 28 Feb 2026

@@ Abstract
Diffusion-based real-world image super-resolution (Real-ISR) methods have demonstrated impressive performance. To achieve efficient Real-ISR, many works employ Variational Score Distillation (VSD) to distill pre-trained stable-diffusion (SD) model for one-step SR with a fixed timestep. However, since SD will perform different generative priors at different timesteps, a fixed timestep is difficult for these methods to fully leverage the generative priors in SD, leading to suboptimal performance. To address this, we propose a Time-Aware one-step Diffusion Network for Real-ISR (TADSR). We first introduce a Time-Aware VAE Encoder, which projects the same image into different latent features based on timesteps. Through joint dynamic variation of timesteps and latent features, the student model can better align with the input pattern distribution of the pre-trained SD, thereby enabling more effective utilization of SD’s generative capabilities. To better activate the generative prior of SD at different timesteps, we propose a Time-Aware VSD loss that bridges the timesteps of the student model and those of the teacher model, thereby producing more consistent generative prior guidance conditioned on timesteps. Additionally, though utilizing the generative prior in SD at different timesteps, our method can naturally achieve controllable trade-offs between fidelity and realism by changing the timestep. Experimental results demonstrate that our method achieves both state-of-the-art performance and controllable SR results with only a single step. The source codes are released at https://github.com/zty557/TADSR
1 Introduction
@@
Figure 1: (a) Comparison between our TADSR(Ours) and PisaSR [21]. In PisaSR, increasing the semantic weight $\lambda_{sem}$ leads to restore more realistic images. As the timestep condition $t$ increases, our model recovers a more realistic parrot image. In contrast, PisaSR shows only an increase in sharpness as $\lambda_{sem}$ increases. (b) The input image and the corresponding outputs of the SD at different timesteps $t$. The outputs vary significantly across different timesteps, reflecting distinct generative priors.
Real-World Image Super-Resolution (Real-ISR) aims to restore high-quality (HQ) images from low-quality (LQ) inputs degraded by complex and unknown factors in real-world scenarios, which has recently attracted increasing attention [25, 35, 40, 15, 9, 38]. To address a broader spectrum of degradation types and achieve more realistic results, many researchers have turned to generative models, particularly diffusion models [10]. Consequently, several works have explored leveraging the generative priors in pre-trained Stable Diffusion (SD) models [20] to tackle Real-ISR, yielding impressive results [24, 18, 22, 33, 9, 34]. Nevertheless, the iterative denoising process inherent in diffusion models introduces significant computational overhead and latency.
@@
To overcome these limitations, some researchers have focused on distilling SD into an efficient one-step model for Real-ISR [32, 21, 3, 8, 39, 37]. Specifically, OSEDiff [32] first leverages the Variational Score Distillation (VSD) loss [27] to distill the SD model, enabling realistic image reconstruction in a single step. Subsequently, S3Diff [39], PisaSR [21], AdcSR [3], and TSDSR [8] also adopt distillation-based approaches to develop SD-based one-step Real-ISR models, using either adversarial loss or modified VSD loss. They typically use a pre-trained SD with trainable LoRA modules as the student model to perform Real-ISR, while employing a fixed-weight SD as the teacher to provide generative guidance.
@@
However, these methods generally fix the timestep (e.g. step 999) injected into the student model while randomly sampling the timestep injected into the teacher model, which prevents them from effectively leveraging the generative prior in SD. Specifically, as shown in Figure 1(b), when the timestep $t$ equals 100, most of the image information is preserved, and the teacher’s output only differs in texture details. As the $t$ grows to 300, the teacher model increasingly activates leaf-related generative priors to predict content that has been obscured by noise. However, with $t$ increasing further to 600, most of the image information is lost, and the teacher model can only recover the overall structure and color of the leaves. This observation indicates that the pretrained SD model exhibits different generative priors at different timesteps. Therefore, existing methods generally suffer from the following two problems: (1) The fixed timestep injected into the student model fails to fully leverage the generative priors at different timesteps in the pretrained SD model; (2) The randomly sampled timestep injected into the teacher model makes it difficult to provide consistent generative guidance. As a result, as shown in Figure 1(a), although we increase the semantic weight $\lambda_{sem}$ in PisaSR [21], it only produces enhanced sharpness without significantly enriching the semantic content. In contrast, our method gradually generates a more realistic parrot as the timestep increases.
@@
In the end, we propose Time-Aware One Step Diffusion Network for Super-Resolution (TADSR), a framework that more effectively distills the generative prior of SD at different timesteps into a one-step diffusion model for Real-ISR. To address the first limitation and better exploit the generative priors at different timesteps, two conditions need to be satisfied: (1) the student model should receive randomly sampled timesteps; (2) the latent features fed into the student model should vary with the timestep, reflecting the noise-level changes in SD. Therefore, we incorporate a Time-Aware VAE Encoder (TAE), which introduces a time embedding layer into the VAE encoder to map the same image to different latent representations based on the timestep. To address the second limitation and ensure consistent generative guidance, we propose a Time-Aware Variational Score Distillation (TAVSD) Loss, which associates the timestep injected into the student model with the one used in the teacher model through a mapping function. When the student model is conditioned on a larger timestep, the teacher receives a latent image corrupted with stronger noise, providing guidance that emphasizes stronger semantic generation in the reconstruction results. Conversely, a smaller timestep leads to similar results with reconstruction, primarily enhancing texture details. Therefore, TAVSD can provide a more consistent generative guidance condition on the injected timestep in the student model. Our contributions are summarized as follows:
  • We propose TADSR, a Time-Aware One-Step Diffusion Network for Real-ISR, which naturally leverages the generative priors of SD at different timesteps to achieve controllable trade-offs between fidelity and realism in Real-ISR. Our TADSR achieves superior performance compared with other SD-based Real-ISR methods on both real-world and synthetic datasets.
  • We propose a Time-Aware VAE Encoder (TAE), which maps the same image to different latent representations based on the timestep, enabling the student model to fully exploit the generative priors at various timesteps.
  • We propose a Time-Aware Variational Score Distillation (TAVSD) Loss, which aligns the timestep of the student and teacher models, providing consistent generative guidance at different timesteps.
2 Related Work
2.1 Real-World Image Super-Resolution
@@
Traditional Image Super-Resolution (ISR) methods [17, 42, 5, 16, 6, 7, 43] typically degrade HQ images using simple downsampling operations to construct HQ-LQ image pairs for training. However, these approaches struggle to handle images degraded by complex real-world processes. To better simulate the unknown and complex degradations in real-world scenarios, several studies [40, 25] have proposed more sophisticated degradation pipelines to synthesize LQ data. Specifically, BSRGAN [40] introduces a random combination of basic degradation operations (e.g., downsampling, blurring, noise) injection, with varying intensities to generate realistic HQ-LQ pairs. Real-ESRGAN [25] proposes a second-order degradation scheme to cover a broader range of degradation types. In addition, inspired by Generative Adversarial Networks (GANs), researchers have adopted adversarial losses to encourage the reconstruction of more realistic images. Although these GAN-based methods can produce richer texture details compared to traditional approaches, they are often unstable to train and prone to generating unnatural artifacts [32].
@@
2.2 Diffusion-Based Real-ISR
Recently, many researchers have leveraged the powerful generative priors of pre-trained diffusion models for Real-ISR tasks to achieve realistic image reconstruction. For example, StableSR [24] conditions the diffusion process on LQ images by injecting them through a learnable time-aware encoder into the SD model, enabling strong detail generation capabilities. DiffBIR [18] utilizes ControlNet to extract structural information from LQ images to better guide the generative prior of SD for SR. PASD [22] and SeeSR [33] extract semantic information from LQ inputs and inject it into SD, resulting in more realistic outputs. Although these approaches yield impressive results, the multi-step denoising process leads to high computational and time costs. To accelerate diffusion-based Real-ISR, OSEDiff introduces the VSD loss to distill the pre-trained SD model, enabling realistic image reconstruction in a single step. S3Diff further adopts degradation-guided LoRA adapters combined with adversarial training to achieve one-step SR. PisaSR trains two separate LoRA adapters for pixel-level and semantic-level guidance, allowing controllable trade-offs between realism and fidelity.
However, these methods overlook the varying generative capabilities of SD in different timesteps and train with a fixed timestep. Our work aims to fully exploit these time-dependent generative prior to achieve superior SR performance and a natural balance between fidelity and realism.
3 Methodology
3.1 Problem Definition
Real-ISR aims to reconstruct HQ images $x_{H}$ from LQ images $x_{L}$ that suffer from complex and unknown degradations. With the advancement of deep learning, researchers have commonly adopted neural networks $G_{\theta}$ to estimate the HQ images and optimize the network through loss functions. The general form of the loss function is:
@@ 3.1 Problem Definition
Recently, with the advancement of diffusion models, several studies [32, 21] have leveraged the generative prior in pre-trained SD and adopted VSD as a regression objective. VSD is designed to align the distribution of generated (fake) images with that of real images. Specifically, a pretrained teacher model trained on real images is used to estimate the score of the real image distribution. In addition, a diffusion model trained on the generator’s outputs provides an estimate of the score of the fake distribution. The generator is then optimized by minimizing the discrepancy between these two score estimates, thereby encouraging it to produce samples that are indistinguishable from real images. The formation of VSD is:
@@
where $\epsilon_{\psi}$ is the pre-trained diffusion model (teacher model) to estimate the real score, $\epsilon_{\phi}$ represents its replica with trainable LoRA (LoRA model) to predict the fake score, and $c$ is a text embedding of a caption describing the input image. $\hat{z}=G_{\theta}(x_{L})$ is the output of the student network $G_{\theta}$, and $\hat{z}_{t}=\alpha_{t}\hat{z}+\beta_{t}\epsilon$ is the noisy latent input. $\epsilon$ is the gaussian noise, and $\alpha_{t}$ and $\beta_{t}$ are the scale parameters in diffusion.
Formally, the VSD loss can be viewed as the residual between the noise outputs of the teacher model and the LoRA model, which can be transformed to the residual of their predicted latent images through $z=\frac{z_{t}-\beta_{t}\epsilon}{\alpha_{t}}$. In practice, the gradient of the VSD loss is applied only to the generator and does not propagate through the teacher model. Therefore, the loss can be expressed in the following form:
$\begin{gathered}\mathcal{L}_{VSD}(\theta)=\mathbb{E}_{t,\epsilon}[\omega^{\prime}(t)\|G_{\theta}(x_{L})-G_{\theta^{-}}(x_{L})\\ -z_{\psi^{-}}(\hat{z}_{t};t,c)+z_{\phi^{-}}(\hat{z}_{t};t,c)\|^{2}_{2}],\end{gathered}$ (3)
where - denotes $stopgrad()$, $z_{\psi}$ and $z_{\phi}$ denote the predicted latent image of the teacher model and LoRA model, respectively. Therefore, we can decode the latent images predicted by the teacher model and the LoRA model into the image space to analyze the guidance provided by VSD.
3.2 Overview
As illustrated in Figure 2, we distill a Student Model $G_{\theta}$ to perform one-step Real-ISR, which consists of a trainable Time-Aware Encoder $E_{\theta}$ and a LoRA finetuned diffusion UNet $F_{\theta}$. Following the OSEDiff [32], we use a pretrained SD model as the Teacher Model $\epsilon_{\psi}$ and its replica with trainable LoRA as the LoRA Model to obtain the fake score. First, we sample an HQ-LQ image pair from the dataset and a timestep $t_{s}$ from a uniform distribution in the range of 0 to 999. Then, both the LQ image and $t_{s}$ are fed into Student Model $G_{\theta}$ to obtain the latent output $\hat{z}_{0}$. We decode $\hat{z}_{0}$ into pixel space and compute the reconstruction loss with the HQ image. In the latent space, we map the $t_{s}$ to another timestep $t_{v}$, and add noise corresponding to timestep $t_{v}$ to the $\hat{z}_{0}$ to obtain $\hat{z}_{t_{v}}$. Then, we feed $\hat{z}_{t_{v}}$ and $t_{v}$ into both the teacher model and the LoRA model to compute the Time-Aware Variational Score Distillation loss $\mathcal{L}_{TAVSD}$ to enhance the realism. Consistent with Figure 1(b), when $t_{v}$ is small, the gradients produced by the TAVSD loss are relatively small and mainly reflect texture details. In contrast, when $t_{v}$ is large, the gradients become significantly larger and provide more semantic guidance. Therefore, the TAVSD loss can offer more consistent gradient guidance condition on $t_{s}$, enabling better distillation of the teacher model. In addition, the LoRA model is trained with a diffusion loss using the data generated by the student model.
@@ 3.2 Overview
3.3 Time-Aware VAE Encoder
In the original SD, as the timestep increases, the input latent features are injected with more noise, thereby activating different generative priors to produce the image. Considering the importance of the timestep in the diffusion process, multi-step SD-based Real-ISR methods often take the timestep as a condition along with the LR image to control the SR process — as seen in the time-aware encoder of StableSR and the ControlNet used in SeeSR and DiffBIR.
However, since multi-step denoising iterations are not required, existing one-step diffusion-based Real-ISR methods generally overlook the role of the timestep and train with only a fixed timestep, making it difficult to fully exploit the generative priors in SD at different timesteps. A straightforward improvement is to randomly sample timestep during training. However, since the original VAE encoder maps the same image to a single latent distribution regardless of the timestep, it is difficult for the diffusion network to effectively activate different generative priors based solely on the timestep, given the same latent input. In the original diffusion process, variations in the timestep reflect changes in the noise level of the latent distribution. We argue that a one-step SD-based Real-ISR network should exhibit a similar property in order to fully leverage the corresponding generative priors. Clearly, directly injecting noise into the latent distribution according to the timestep is inappropriate, as it would compromise reconstruction fidelity.
Therefore, we propose a Time-Aware VAE Encoder (TAE) to better utilize the generative priors in SD. By incorporating a temporal embedding layer into the VAE encoder, TAE encodes the input image into different latent distributions conditioned on the timestep $t_{s}$, enabling synchronized variation between $t_{s}$ and latent distribution, thus better activating the generative priors at different timesteps within SD. This process can be formulated as: